SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Friday 27 June 2003

Status Report From: Goddard Space Flight Center
Posted: Friday, June 27, 2003

ANNOUNCEMENTS

* We expect to lose the high gain antenna link today (or at least not be able to reacquire it on tomorrow's 34 meter pass). We no longer have it on 26m stations, just on the 34m stations. Once we lose the high gain antenna link on the 34m stations, the spacecraft will be transitioned to low rate.

We have daily passes on 34 meter stations for monitoring and commanding of the spacecraft. The 34 meter schedule for this week is included below. Since the times are very limited, we'd like to request that ALL TEAMS BE AVAILABLE (at EOF or by phone) during the passes, in case your instrument comes up in an unexpected state. When downlink only 34 meter stations are scheduled, a 26 meter station will be used for uplink (i.e. all 34 meter passes will have commanding capability).

Week 26 34-Meter Passes

                   BOT   EOT
                 --------------
   Fri     DSS-54 13:00-17:45 UT
   Sat     DSS-24 17:35-20:35 UT
   Sun     DSS-54 13:00-17:30 UT

* Current RHESSI default target is AR10391. When selecting this target to maximize coordination w/others, use campaign number 6850.

* There will be about 10 UMD physics engineering students visiting the EOF through the Goddard Outreach Program today at 11:15 AM. POC: Bob Dutilly.

FOT REPORT

Spacecraft Status: Nominal, except for the HGA pointing mechanism

HGA Position Monitoring Disabled

HGA Nominal Control Disabled

Spacecraft Anomalies: None

Accomplished Activities: SWAN (x2)

Planned Activities: RSL, New Tracking Star, SWAN

Upcoming Operations: None

Ground Anomalies: 178/0125 TCP unable to lock on TM signal. D46. HGA offpointing. SNRs ~2.5. Too low for station to lock.

SOLAR STATUS

Web Page for Planning: http://www.bbso.njit.edu/arm

Magnetic Maps and Heliospheric Forecasts: http://www.lmsal.com/forecast/

INSTRUMENT STATUS

CDS: Safed - doors closed.

UVCS: Safed - door closed.

LASCO: Safed - doors closed.

EIT: Bakeout.

MDI: Nominal. Will safe when we get to low rate.

TRACE: Nominal. Continues multi-wavelength observations of AR10391 and AR10387.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter